OpenStudy (anonymous):

Simplify rad (96) + 4 rad (27) - rad (150) + rad (75)

OpenStudy (saifoo.khan):

\[\huge -\sqrt{6}+17\sqrt{3}\]

OpenStudy (anonymous):

are you looking for a solution to: \[\sqrt{96} + 4\sqrt{27} -\sqrt{150} + \sqrt{75}\] ?

OpenStudy (anonymous):

yess

OpenStudy (saifoo.khan):

My answer's right.

OpenStudy (anonymous):

to break down a square root like \[\sqrt{96}\] You just start dividing 96 by the sqare numbers (e.g 4, 9, 16, 25, 36...). Once you find the biggest one that goes in you can split it up: \[\sqrt{96} = \sqrt{3 \times 36} = \sqrt{3} \times \sqrt{36} = \sqrt{3} \times 6\]